# Can Language Models Reason about

### v1

@@ 3 Can LMs Reason about Individualistic Human Values?
How well can LMs reason about an individual’s values after observing value-expressing statements from that same individual? Figure 3 shows the evaluation of various LMs’ ability to reason about individualistic values. All models substantially outperform the Random baseline, where a statement is chosen randomly from each question group. The GPT-4o (0806) Rand baseline,in which GPT-4o is given no demonstrations, achieves higher accuracy than Random, suggesting that GPT-4o has systematic preferences over statements, allowing it to align with broader human preferences even without demonstrations. Notably, GPT-4o with 200 demonstrations considerably outperforms the model without demonstrations (63.5 vs. 54.8), indicating that individual value demonstrations can effectively guide LMs in interpreting their general value preferences. Yet, no model achieves particularly high performance on the task, with average performance only ranging between 55% to 65%. Lastly, certain categories of statements (e.g., Happiness & Well-being, Ethical Values & Norms) are easier to predict than others (e.g., Economic Values, Postmaterial Index). Please refer to Figure 8 in §B.2 for how each type of value statements influences the prediction of other types.
@@
Whose values are easier for LMs to predict? As shown in Figure 4 (blue boxes), LMs exhibit uneven performance across demographic groups, indicating varying difficulty levels in predicting values across populations. For instance, Llama-3.1-8B is most accurate at predicting values for individuals from Oceania, with high income, and from the upper-middle-class. These disparities across sub-populations align with findings from prior research that probed LMs using general multiple-choice questions from the WVS, comparing the model’s output distributions to human labels (Durmus et al., 2024). Refer to Figure 9 in Appendix §3 for full results showing performance disparity across other demographics groups for GPT-4o, and Figure 11 to 21 for Llama-3.1-8B.
@@
Figure 4: Comparing Llama-3.1-8B zero-shot vs. IndieValueReasoner performances broken down by demographics groups across the Continent, Income, and Social Class demographics dimensions. The lower the $\sigma$, the more impartial the performance of the IndieValueReasoner is in reasoning about individualistic values across populations with different demographic groups.
@@
How impartial or equitable are LMs in their reasoning about individuals across demographics? Table 3 shows the Value Inequity Index ($\sigma$Inequity) of various frontier LMs. Notably, models with similar proficiency in individualistic value reasoning (indicated by accuracies in Figure 3) may have drastically different $\sigma$Inequity, revealing discrepant equity levels regarding diverse populations. For instance, both GPT-4o (0513) and Llama-3.1-70B have an accuracy of 63.7, showing a similar proficiency level. However, GPT-4o (0513) has higher $\sigma$Inequity (2.87), compared to Llama-3.1-70B (1.94), indicating a less equitable value representation. We introduce $\sigma$Inequity as a new quantifiable measure of the impartiality or equity of LMs. $\sigma$Inequity presents complementary metrics to proficiency for assessing LMs’ capability for reasoning about individualistic human values and achieving the potential of building models for all.
@@
How does the number of demonstration statements impact model’s predictions? Figure 5 shows the results of evaluating the impact of varying the number of demonstration value-expressing statements. As expected, including more demonstration statements leads to higher accuracy for GPT-4o. However, it’s noteworthy that even with as few as 50 demonstration examples, the model’s accuracy improves from 54.79 to 60.59, demonstrating the effectiveness of a relatively small number of examples in guiding the model to grasp individual values.
How informative is general demographics information for LMs in predicting individualistic value choices? Figure 5 compares probing setups with and without demographic information. When only demographic data is provided (leftmost orange box), GPT-4o achieves a performance score of 60.31, slightly lower than 60.59 when 50 value-expressing statements are included. Combining a varied number of value statements with demographic information consistently results in marginally higher performance compared to setups without demographic information, although the difference is not statistically significant GPT-4o. Notably, when the model is given more value-expressing statements, it achieves higher accuracy than when provided with fewer statements alongside demographic information. This suggests that value statements capture significant latent information about individualistic values, necessary for approximating the uniqueness of individuals. For weaker models like GPT-4o-mini, including demographics leads to significantly better predictions compared to providing value statements alone as the model likely has more difficulty in interpreting descriptive value statements (see more details in Figure 10 in §B.2). Importantly, relying solely on demographic information to infer individual values may inadvertently reinforce stereotypical group-based interpretations, undermining a nuanced and precise understanding of individual values.
4 How does training models on people’s value expressions reveal patterns and dynamics of individualistic values?
@@ 4.1 Method
[…]
Our goal in training the IndieValueReasoner is not to “solve” the individualistic value reasoning mission, but rather to conduct a deeper examination of how data and LMs can be combined to uncover meaningful patterns in human values and to assess the data-driven upper-limit performance for this task. To show the comparative trend, we include both statistics-based and LM-based baselines. For statistics-based methods, we consider selecting the statement for $I_{i}\in\mathbb{I}_{\text{eval}}$ based on (1) Global (majority vote): the majority vote across the global pool of individuals ($\mathbb{I}_{\text{train}}$); (2) Resemble (top 1): the statement choice of $I_{j}\in\mathbb{I}_{\text{train}}$ who shares the most number of common demonstration statements with $I_{i}$; (3) Resemble (top cluster): the majority vote among the top cluster of training individuals who share the most number of common demonstration statements with $I_{i}$. For LM-based baselines, we consider (1) GPT-4o (no demo.): GPT-4o without demonstrations; (2) GPT-4o (only demographics): GPT-4o with only demographics information; (3) GPT-4o (200 demo.): GPT-4o with 200 demonstrations; (4) Llama-3.1-8B (200 demo.): Llama-3.1-8B with 200 demonstrations. Baselines details are shown in §C.1. Finally, we evaluate models with both polar and refined evaluation setups, where we present demonstration and probe statements in those two granularity forms, respectively.
@@
Method Probe 1 Probe 2 Probe 3 Avg. Probe 1 Probe 2 Probe 3 Avg. Avg.
Random 46.37 45.51 44.23 45.37 29.16 29.03 25.43 27.87 36.62
Resemble (top 1) 70.31 70.15 69.02 69.83 53.26 54.01 53.27 53.51 61.67
Resemble (top cluster) 74.74 74.87 71.60 73.73 59.32 60.78 58.32 59.47 66.60
@@ 4.2 Results
Training LMs with individualistic value statements results in proficient IndieValueReasoners. Table 4 shows the accuracy of various IndieValueReasoner models compared to baselines with both polar and refined evaluation sets. [probe=p+r,demo=mix:200,N=1600], the best-performing IndieValueReasoner model achieves 46.6% of relative improvements compared to the zero-shot setting, [Llama-3.1-8B (200 demo.)]. Compared to [GPT-4o (only demographics)], the best-performing GPT-4o configuration, [probe=p+r,demo=mix:200,N=1600] achieves 34.0% of relative improvement, showing that the smaller and less capable models can substantially improve over larger models with supervision of individualistic values data. Moreover, we compare IndieValueReasoner trained to predict both refined and polar question forms. The model solely trained to select among coarse statement options, i.e., [probe=p,demo=mixed,N=800], does well only on polar test cases without extrapolating to refined statements. The model solely trained on refined statements, i.e., [probe=r,demo=mixed,N=800], improves on refined test cases, while maintaining performance on polar questions, despite not being as high as the model specialized in polar questions. We choose to combine both refined and polar probes for training to have a balanced performance between the two forms. We further show that training data with a mixed number of demonstrations, i.e., [probe=p+r,demo=mixed,N=800], achieves better performance (66.49) compared to the model trained with a fixed number of 200 demonstration statements (66.14), [probe=p+r,demo=200,N=800], when both are tested against examples with 200 demonstrations. This shows that despite we seemingly provide less information during training (i.e., less total number of demonstration statements for [probe=p+r,demo=mixed,N=800]), the diversity brought by the mixed number of demonstrations provides richer variety of information for the model to gain stronger generalizability. Even better, combining data with a both 200 and a mixed number of demonstrations results in the best-performing model, [probe=p+r,demo=mixed+200,N=800]. Finally, we further scale up the training data size by including 1.6K individuals’ data per probe, [probe=p+r,demo=mixed+200,N=1600], further improving the performance. Specifically, Figure 6 (Left) shows that the increased training data size consistently results in improved performance of IndieValueReasoner when tested with different numbers of demonstration statements, highlighting the importance of data scale.
Individuals with similar value demonstration trajectories are informative for predicting a new individual’s value choices. Statistics-based baselines all have Oracle access to the data of all individuals. Searching and aggregating value choices of similar individuals offers strong predictive power in facing the value choices of new individuals, especially when we aggregate opinions of a cluster of individuals with similar value judgment trajectories, as shown by [Resemble (top cluster)]. These statistics-based baselines all substantially outperform all zero-shot LM-based baselines. This result highlights that off-the-shelf LMs risk guessing individual value choices without explicitly teaching. However, notably, [probe=p+r,demo=mix:200,N=1600], the best-performing IndieValueReasoner (67.67) beats [Resemble (top cluster)] (66.60) despite it has only seen demonstration sequences from 1.6K individuals per question, rather than the entirety of 92K individuals as for statistics-based baselines. This shows a relative sample efficiency and stronger generalizability of employing LMs for capturing individual value patterns.
IndieValueReasoner has improved $\sigma$Inequity compared to zero-shot LMs, highlighting the importance of teaching individual differences for equitable models. In addition to the improved reasoning proficiency, [probe=p+r,demo=mix:200,N=1600] also achieves improved $\sigma$Inequity (2.22) compared to zero-shot Llama-3.1-8B (2.97). Specifically, Figure 4 shows a breakdown view of how the individualistic value reasoning ability increases more in previously under-performed demographics groups, For instance, IndieValueReasoner has +18.24% absolute performance gain among individuals from the lowest-performing region, South America, more than the better-performing regions like North America (+14.38%) and Oceania (+12.51%). This shows that training models on extensive global individuals’ data helps alleviate the partiality of off-the-shelf Llama-3.1-8B in reasoning about individual differences across demographic groups. Breakdowns of all demographics dimensions are shown in Figure 11-21 and Table 19 in §C.2.
@@
A hybrid number of demonstrations improves reasoning generalizability. As shown in Figure 6 (Right), across all models, increasing the number of test demonstrations improves the model’s ability to infer an individual’s value choices. Interestingly, training IndieValueReasoner with a randomized mix of demonstrations (between 50 to 200) results in a better performance than training with any fixed number of statements. Counterintuitively, using the maximum number of demonstrations (200) only produces a moderately effective model, even when tested in the same 200-demonstration format. This model performs poorly when fewer demonstrations are given during testing, where stronger extrapolation abilities are needed to make accurate inferences based on limited evidence. Conversely, a model trained on fewer demonstrations (50) excels at making inferences with little evidence but struggles to generalize when given more specific demonstrations. Training on a randomized number of demonstrations (50 to 200) performs well, except when tested with 150 or 200 demonstration statements. To address this gap, we developed a hybrid model, trained on both a randomized number of demonstrations and the full sequence of 200 demonstration statements, showing the best performance. These results demonstrate the synergistic power between data with different demonstration configurations for improving the individualistic value reasoning to generalize with both abstract and specific evidence.
@@
How do models trained on data from different global regions show discrepant predictive power over cross-region individuals? In order to gauge how individual data across different global regions impact a learned model’s ability to reason about the value of diverse populations, we train models with data from each of the six continents (see Figure 7). Indeed, content-specific models result in drastically divergent continent-specific performances. These models typically achieve the best (Europe, North America, Asia) or second-best (South America, Africa) performance for the corresponding continent’s test population (except Oceania), highlighting the strong influence of regional data in supervising performance on the same population. Sometimes, we also observe a particularly strong performance of some content-specific models on other populations. For instance, North America model achieves the best performance on the South America test data; European model achieves the best performance on Africa test set. This trend aligns with geographical proximity and the commonly held impression of a close influence between the source and test continents. The Oceania model and performance on the Oceania test sets prove exceptional cases among continent-specific models: all models (except the Africa model) show quite high performance on the Oceania test set, and the Oceania model performs poorly across all continent of test sets, except on its own population and the North American population, which shares cultural similarity. We hypothesize that such an irregular pattern is due to the Oceania data lacking diversity, as all Oceania data is from New Zealand. Thus, a model trained on a relatively homogeneous pool of individuals cannot learn the diverse individualistic value patterns; correspondingly, a homogeneous test set is easier to predict, even for regional models. Finally, the model trained on worldwide data achieves comparable, if not stronger, performance on all continent-specific test sets compared to regional models. These results highlight the importance of diverse cross-region data for teaching the models a robust sense of global human value patterns.
@@
Table 5: Models trained with only value-expressing statements, with only demographics descriptions, or both.
@@
Training model solely on demographics descriptions of individuals does poorly in test cases with descriptive value-expressing statements. We experiment with training a IndieValueReasoner using only demographics descriptions (e.g., “I’m 25-34 years old …”) instead of descriptive value-expressing statements. Such a model cannot learn to generalize to test cases with descriptive value-expressing statements as demonstration examples. Similarly, the model trained from descriptive value statements also struggles to make predictions based on demographics demonstrations (though with a better overall performance). Surprisingly, training with a combination of demographic-based demonstrations and value-expressing statements improves performance in both test scenarios, outperforming models trained on either data type alone. This suggests a mutually reinforcing effect between demographic information and value-expressing statements.
5 Discussion
5.1 Limitations and Future Directions

### v2

@@ 3 Can LMs Reason about Individualistic Human Values and Preferences?
Figure 2 shows that all models outperform the Random baseline, where a statement is chosen randomly. The GPT-4o (0806) Rand baseline, in which GPT-4o is given no demonstration, achieves higher accuracy than Random, suggesting that GPT-4o has systematic preferences over statements, allowing it to align with broader preferences without demonstrations. Notably, GPT-4o with 200 demonstrations considerably outperforms the model without any (63.5 vs. 54.8), indicating that demonstrations can effectively guide LMs in interpreting an individual’s general value preferences. Yet, no model achieves high performance, with average accuracies ranging between 55% to 65%. Lastly, certain categories of statements (e.g., Happiness & Well-being, Ethical Values & Norms) are easier to predict than others (e.g., Economic Values, Postmaterial Index). Figure 8 in Appendix §C.2 shows how each value statements type influences the prediction of others.
Whose values are easier for LMs to predict?
@@
LMs are subject to varying difficulty levels in predicting values across populations, e.g., Llama-3.1-8B is most accurate at predicting values of individuals from Oceania, as shown in Figure 3 (blue boxes). The disparity across sub-populations aligns with prior research that compares models’ distributional outputs to human labels in WVS (Durmus et al., 2024). Refer to Figure 9 in Appendix §3 for full results of other demographics groups for GPT-4o, and Figure 11 to 21 for Llama-3.1-8B.
@@
Figure 3: Comparing Llama-3.1-8B zero-shot vs. IndieValueReasoner performances broken down by Continent. Lower $\sigma$ indicates more impartial model.
Equity in individualistic value reasoning.
@@
Table 2 shows the Value Inequity Index ($\sigma$Inequity) of various frontier LMs. Notably, models with similar proficiency in individualistic value reasoning (indicated by accuracies in Figure 2) may have drastically different $\sigma$Inequity, revealing discrepant equity levels across populations. For instance, both GPT-4o (0513) and Llama-3.1-70B have an accuracy of 63.7. However, GPT-4o (0513) has higher $\sigma$Inequity (2.87), compared to Llama-3.1-70B (1.94), indicating a less equitable value representation. We introduce $\sigma$Inequity as a new quantifiable measure of the equity of LMs. $\sigma$Inequity presents complementary metrics to proficiency for assessing LMs in individualistic value reasoning.
@@
Impact of the number of demonstration statements.
Figure 4 shows that increasing the number of demonstration statements improves GPT-4o’s accuracy. Notably, even 50 statements boost accuracy from 54.79 to 60.59, demonstrating the impact of a small set of value expressions in guiding models to understand individualistic values.
Demographics descriptions do not teach LMs accurate sense of individualistic values.
Figure 4 shows that when only demographic statements are provided (leftmost orange box), GPT-4o achieves 60.31% accuracy, slightly lower than the 60.59% achieved with 50 value-expressing statements. Adding demographic statements alongside value statements consistently yields marginally higher performance, though not statistically significant. Notably, when GPT-4o is given more value demonstrations, its accuracy improves compared to scenarios with fewer value statements and demographic data. This indicates that value statements encapsulate key latent information essential for approximating individual uniqueness. For weaker models like GPT-4o-mini, including demographics significantly improves predictions compared to value statements alone, likely due to challenges in interpreting descriptive value expressions (see Figure 10 in Appendix C.2). Importantly, relying solely on demographics risks reinforcing stereotypical group-based interpretations, undermining nuanced understanding of individual values.
Please refer to §C.2 for full probing experiments.
4 Training Models to Capture Patterns and Dynamics of Individualistic Values
@@ 4.1 Method
The rich data in IndieValueCatalog allows us to train a series of Individualistic Value Reasoner models (IndieValueReasoner) based on Llama-3.1-8B. We form the training data using value statements from $\mathbb{I}_{\text{train}}$. Each training data contains $d$ demonstration statements ($d=$200 or mixed stands for drawing 200 or randomly between 50-200 demonstrations, respectively. and statement candidates of a probing question (probe), all from the same individual. The model takes in the demo statements and outputs a choice among the probe candidates. Both demo and probe can take either polar (p) or refined (r) forms. For each of the 253 questions ($q$), we sample N individuals from $\mathbb{I}_{\text{train}}$ to form different demonstration sets for $q$, and use each individual’s statement choice of $q$ as the gold label, forming 253×253$\times$253 ×N training data. Full training details are shown in Appendix §D.1.
Our goal in training the IndieValueReasoner is not to solve the individualistic value reasoning mission but to explore how data and LMs can be combined to reveal meaningful patterns in human values and assess the data-driven performance upper-bound for this task. We include both statistics and LM-based baselines. For statistics-based baselines, we consider selecting the statement for $I_{i}\in\mathbb{I}_{\text{eval}}$ based on (1) Global (majority vote): the majority vote across the global pool of individuals ($\mathbb{I}_{\text{train}}$); (2) Resemble (1600/all, top 1): the statement choice of $I_{j}\in\mathbb{I}_{\text{train}}$ who shares the most number of common demonstration statements with $I_{i}$, among 1,600 randomly selected or the overall pool of individuals; (3) Resemble (1600/all, top cluster): the majority vote among the top cluster of training individuals who share the most number of common demonstration statements with $I_{i}$, among 1,600 randomly selected or the overall pool of individuals. For LM-based baselines, we consider (1) GPT-4o (no demo.): GPT-4o without demonstrations; (2) GPT-4o (only demographics): GPT-4o with only demographics information; (3) GPT-4o (200 demo.): GPT-4o with 200 demonstrations; (4) Llama-3.1-8B (200 demo.): Llama-3.1-8B with 200 demonstrations. Details of baseline are shown in Appendix §D.1. Finally, we evaluate models using both polar and refined setups for both demonstrations and probes.
@@
Random 45.37 27.87 36.62
Global (majority vote) 64.95 48.70 56.83
Resemble (1600, top 1) 65.65 47.55 56.60
Resemble (1600, top cluster) 70.39 54.36 62.38
Resemble (all, top 1) 69.83 53.51 61.67
Resemble (all, top cluster) 73.73 59.47 66.60
GPT-4o (no demo.) 54.79 33.06 43.93
GPT-4o (only demographics) 60.31 40.69 50.50
GPT-4o (200 demo.) 63.46 35.59 49.52
Llama-3.1-8B (200 demo.) 54.34 37.97 46.16
[probe=p,demo=mixed,N=800] 73.59 44.08 58.84
[probe=r,demo=mixed,N=800] 73.25 59.94 66.59
[probe=p+r,demo=200,N=800] 73.45 58.84 66.14
[probe=p+r,demo=mixed,N=800] 73.59 58.27 66.49
[probe=p+r,demo=mixed+200,N=800] 74.29 60.23 67.26
[probe=p+r,demo=mixed+200,N=1600] 74.74 60.60 67.67
Table 3: Results of IndieValueReasoner models for improved individualistic value reasoning for both the polar and refined evaluation setups. For the middle section of ablation models, the best performances are bolded, and the second best performances are underlined. All results in this table are obtained by giving 200 demonstration value-expressing statements during test time. For naming convention, [probe=p+r,demo=mixed+200,N=800] refers to a model trained using both mixed and fixed 200 demonstration statements, and evaluated with probing statements in both polar and refined forms. Each of the 253 value questions has 200 examples for each setup—(mixed, refined), (mixed, polar), (200, refined), and (200, polar)—totaling 800 examples. Please refer to Table 17 for details on the baseline configurations.
@@ 4.2 Results
IndieValueReasoners improves over individualistic value reasoning.
Table 3 shows that the best-performing IndieValueReasoner, [probe=p+r,demo=mix:200,N=1600], achieves 46.6% relative improvements compared to the zero-shot setting, [Llama-3.1-8B (200 demo.)]. Compared to the best-performing GPT-4o configuration, [GPT-4o (only demographics)], the best trained model achieves 34.0% relative improvement, showing that smaller and less capable models can greatly improve over more capable models with supervision of individualistic values data.
We compare models trained to predict refined and polar question forms. The polar-specialized model, [probe=p,demo=mixed,N=800], does well only on polar questions without extrapolating to refined questions. The refined-specialized model, [probe=r,demo=mixed,N=800], improves on refined questions, while maintaining performance on polar questions, despite not as high as the polar-specialized model. We choose to combine both refined and polar probes for training to have a balanced performance between the two forms.
We show that using a mixed number of demonstrations ([probe=p+r,demo=mixed,N=800]) improves performance (66.49) over a fixed 200 demonstrations ([probe=p+r,demo=200,N=800]) when tested on 200-demonstration examples (66.14). This shows that despite we seemingly provide less information during training (i.e., less total number of demonstration statements for [probe=p+r,demo=mixed,N=800]), the diversity brought by the mixed number of demonstrations provides richer variety of information for the model to gain stronger generalizability. Even better, combining data with 200 and mixed demonstrations produces the best-performing model, [probe=p+r,demo=mixed+200,N=800].
Finally, scaling up the training data to 1.6K individuals ([probe=p+r, demo=mixed+200, N=1600]) further enhances performance. Figure 5 (Left) shows that increasing the data size consistently improves IndieValueReasoner’s performance across varying numbers of demonstrations.
@@
Similar value demonstration trajectories can help predict an individual’s value choices.
Statistics-based baselines rely on oracle access to data from all individuals, allowing for strong predictive power by aggregating value choices from similar individuals, as demonstrated by [Resemble (top cluster)]. These baselines significantly outperform zero-shot LM-based approaches, which risk guessing individual value choices without explicit teaching. However, the best-performing trained model, [probe=p+r,demo=mix:200,N=1600] (67.67), outperforms [Resemble (top cluster)] (66.60), despite using data from only 1.6K individuals per question—substantially fewer than the 92K used by statistics-based baselines. When compared against a statistics-based baseline using the same 1.6K randomly sampled individuals, the trained model achieves a significantly higher accuracy (67.67 vs. 62.38). This shows the superior sample efficiency and generalizability of LMs in capturing individual value patterns.
Improved $\sigma$Inequity.
The best IndieValueReasoner achieves improved $\sigma$Inequity (2.22) compared to zero-shot Llama-3.1-8B (2.97). Figure 3 highlights the performance improves greatly among previously underperforming demographic groups. For example, IndieValueReasoner achieves an +18.24% absolute performance gain in South America, the lowest-performing region, compared to Oceania (+12.51%) and North America (+14.38%). This shows that training models on diverse, globally representative data alleviates the partiality of off-the-shelf Llama-3.1-8B in reasoning about demographic differences. Breakdowns of all demographic dimensions are shown in Figure 11-21 and Table 19 in Appendix §D.2.
@@
A hybrid number of demonstrations improves reasoning generalizability.
Figure 5 (Right) shows that increasing the number of test demonstrations enhances models’ performances. Notably, training IndieValueReasoner with a random mix of 50 to 200 demonstrations outperforms training with any fixed number of statements. Yet, a model trained with a maximum of 200 demonstrations performs only moderately well but struggles with fewer test demonstrations, where stronger extrapolation is needed. Conversely, a model trained on 50 demonstrations excels with limited evidence but struggles to generalize with more. Training with a random number of demonstrations (50 to 200) yields strong overall performance but underperforms when tested with 150 or 200 demonstrations. Thus, we trained a model on both mixed demonstrations and the full 200, achieving the best results. These findings highlight the synergy between diverse demonstration configurations for improving individualistic value reasoning across both abstract and specific evidence.
@@
Table 4: Models trained with value-expressing statements, demographics descriptions, or both.
Models trained on different global regions show discrepant predictive power over cross-region individuals.
To evaluate how regional data impacts a model’s ability to reason about diverse populations, we trained models using data from each of six continents (Figure 6). These continent-specific models showed significantly varying performances. They typically achieve the best (Europe, North America, Asia) or second-best (South America, Africa) performance for the corresponding continent’s test population (except Oceania), emphasizing the strong impact of regional data on performance for the corresponding populations. Sometimes, we also observe a particularly strong performance of some content-specific models on other populations. For instance, North America model achieves the best performance on the South America test data; European model achieves the best performance on Africa test set. This trend aligns with geographical proximity and the commonly held impression of a close influence between the source and test continents.
Oceania was an exception: while most models (except the Africa model) performed well on Oceania’s test set, the model trained on Oceania data underperformed on all test sets except Oceania and North America. We hypothesize this is due to the limited diversity of the Oceania dataset, which consists entirely of responses from New Zealand.
To quantitatively confirm the homogeneity of the Oceania data, we assess whether respondents’ answers are skewed across survey questions (i.e., disproportionately clustered around certain options). To measure this, we compute the Shannon’s Evenness Index ($J^{\prime}$). This metric captures how evenly distributed responses are across answer choices, analogous to species evenness in ecology. The index ranges from 0 (completely uneven) to 1 (perfectly even), with lower values indicating more homogeneous response patterns.
$J^{\prime}=\frac{H^{\prime}}{H^{\prime}_{max}}=\frac{H^{\prime}}{\ln(S)}$
where $H^{\prime}=-\sum_{i=1}^{S}p_{i}\ln(p_{i})$
Figure 7: The Shannon’s Evenness Index ($J^{\prime}$) across different continents. Lower $J^{\prime}$ values indicate a more uneven distribution across all answer choices.
As we can see from Table 7, Oceania indeed has a lower $J^{\prime}$ score compared to other continents, meaning that it has a more uneven distribution across all answer choices. This exactly confirms that Ocreania data cluster around certain answer choices, thus proving to be less diverse. This show that a homogeneous training set limits the model’s ability to capture diverse value patterns; correspondingly, a homogeneous test set is easier to predict, even for regional models.
Finally, the globally trained model consistently matched or outperformed continent-specific models across all regions, emphasizing the importance of diverse, cross-regional data for robust reasoning about global human value patterns.
Training on demographic descriptions does not generalize to test cases with value-expressing statements.
@@
We train a IndieValueReasoner using only demographic descriptions (e.g., “I’m 25-34 years old”) instead of value-expressing statements. As shown in Table 4, this model fails to generalize to test cases using value-expressing demonstrations. Conversely, models trained on value-expressing statements struggle with demographic-based demonstrations but perform slightly better overall (68.14 vs. 65.62). Training with a mix of demographic and value-expressing data improves in both test cases (70.88), suggesting a mutually reinforcing relationship between the two data types.
5 Discussion and Future Directions
Impact.
Predicting individual value statements in NLP has significant practical value, particularly for applications requiring a deep understanding of human behavior. For instance, mental health chatbots can offer more empathetic and context-aware support, while personalized recommendation systems can improve content relevance by aligning with user values. Moreover, understanding individual values is essential for developing culturally sensitive and inclusive AI, enabling more nuanced human-AI interactions.
Predicting individualistic human values also intersects with existing computational social science research (CSS) that simulates human behavior via LLMs. Recent work explores using LLMs to simulate social interactions and cultural dynamics Ziems et al. (2024); Zhou et al. (2024); Park et al. (2022). Finally, computationally reasoning through individual human values also potentially add novel insights to other disciplines, such as moral philosophy or psychology. By bridging technical innovation into humanity research, value prediction not only refines AI systems for real-world alignment but also offers ethically scalable methods to study decision-making, enriching both technology development and social scientific inquiry.
Future.